# Reevaluating Policy Gradient Methods for Imperfect-Information Games

### v2

@@ 5 Experiments
While our high-performance exploitability computations open the door to evaluating model-free DRL for IIGs more rigorously, the novelty of these benchmarks necessitates particular care in experimental design and interpretation. Simply showing that tuned PG methods perform “well” on these benchmarks would not provide strong evidence for the policy gradient hypothesis, as there are no numbers for FP-, DO-, or CFR-based DRL algorithms against which to compare. The obvious recourse—which we take—is to report our own results for FP, DO, and CFR-based DRL algorithms. However, since we have hypothesized these algorithms will underperform, our role in tuning them should be carefully scrutinized. To aid readers in this endeavor, we provide detailed documentation of our algorithm implementations, hyperparameter tuning procedure, and hyperparameter tuning results, in addition to final results.
@@ 5.1 Algorithms and Implementation
Figure 2: Exploitability results. For each game-algorithm pair, the box and whisker depicts the distribution of final exploitability over the runs from the hyperparameter tuning launch (left) and evaluation launch (right) with square-root y-axis scale. R-NaD, NFSP, ESCHER, and PSRO failed to outperform generic PG methods (MMD, PPO, PPG).
@@ 5.2 Design
For the hyperparameter tuning launch, for each combination of the 4 games and 7 algorithms, we test 50 hyperparameter configurations for 10 million steps over 3 seeds. To create these configurations, we modify the default positive real-valued hyperparameters by independent randomly sampled powers of 2—multiplying for the vast majority of hyperparameters, but exponentiating for some hyperparameters upper bounded by 1. We take the default values for hyperparameters from the implementation source in almost all cases. The most notable exceptions to this are the network architecture and optimizer, for which we impose a 3-layer 512-hidden unit fully connected network and Adam (Kingma and Ba, 2015) on every algorithm, and the entropy coefficient, for which we apply the value from Sokota et al. (2023) to all generic PG methods. We provide a comprehensive list of hyperparameters, default values, and links to the specific lines of code of the external sources from which those default values were taken in Appendix G.
For the evaluation launch, we run the 5 strongest hyperparameter configurations from the first launch, as measured by lowest average final exploitability, with 10 fresh seeds, for 10 million steps, for each combination of the 4 games and 7 algorithms. From these evaluation runs, we report both exploitabilities and head-to-head comparisons. For the head-to-head comparisons, we select the policy with the lower median final exploitability over the 10 seeds for each hyperparameter configuration, as well as an approximate Nash equilibrium policy computed using discounted CFR (Brown and Sandholm, 2019a), and compare these selected policies for each pair of algorithms for each game.
@@ 5.3 Results
We summarize two of the main findings from our experiments: exploitability performance and head-to-head performance.
@@
Figure 3 plots head-to-head comparisons for each pair of algorithms. Relative performance patterns are similar to those for exploitability: The generic PG methods matched or defeated other algorithms—and even generally outperformed the approximate Nash equilibrium against PSRO—while maintaining rough parity among themselves; R-NaD either approached the performance of the generic PG methods or underperformed them, while NFSP and, especially, ESCHER and PSRO were noncompetitive.
@@ 6 Discussion
However, we emphasize that we make no claim—and should not be interpreted as having claimed—that we have established as true the policy gradient hypothesis or any other speculation discussed in this work. The prudent interpretation of our experiments is nuanced and gives due consideration to multiple limitations.
First, the games on which we performed experiments are homogeneous: 1) Imperfect information is introduced via hidden actions; 2) there is no chance; 3) there are only terminal nonzero rewards, which are binary; and 4) actions represent piece placements onto a 3-by-3 board. We do not believe that the relative performance of the algorithms we evaluated would be universally consistent across adversarial IIGs at large (indeed, it is not even consistent across these homogeneous games).
@@ 7 Conclusion
This work hypothesizes and presents evidence that generic deep PG methods are strong approaches for adversarial IIGs. Our results, along with those of Yu et al. (2022), who demonstrated PPO’s effectiveness in cooperative IIGs, and Sokota et al. (2023), who demonstrated MMD’s effectiveness in adversarial IIGs, add to a growing body of evidence for the possibility that a single, simple PG method could serve as a universal algorithm for DRL in games. We believe this possibility is an aspiration worth pursuing.
@@ 8 Acknowledgments
9 Erratum
arXiv version 1 claims to show results for MMD with constant reverse KL coefficients. In fact, it shows results for scheduled reverse KL coefficients that decrease over time. We thank Chang Lei for bringing this to our attention.
@@ Appendix A The exp-a-spiel Package
We release a Python library to efficiently compute exploitability of policies for the four benchmark games introduced in Section 4. The library supports outputting state representations compatible with OpenSpiel [Lanctot et al., 2019], making it straightforward to evaluate the exploitability of policies trained using the latter library.
@@ A.1 Exploitability Computation
Internally, exp-a-spiel evaluates the exploitability of policies by carrying out computations using the so-called sequence-form representation of the four games [Romanovskii, 1962, von Stengel, 1996, Koller et al., 1996]. This representation is roughly 1000 times more compact than the actual game trees, allowing us to compute exact exploitability (and head-to-head values) on commodity hardware. This representation is comprised of two objects: (i) the treeplexes $\cal X,\cal Y$ of the two players [Hoda et al., 2010], and (ii) the sequence-form payoff matrix $A$ of the game. Every policy for player 1 admits an equivalent vector $x\in\cal X$ (resp. $y\in\cal Y$ for player 2), and the expected payoff corresponding to player 1 can be computed in closed form via the bilinear form $x^{\top}Ay$. Due to the size of the games, the sequence-form payoff matrix $A$ is not stored explicitly in memory by exp-a-spiel, but is rather materialized on the fly in a multi-threaded fashion as needed.
@@
    Then, the vectors $g_{1}\coloneqq Ay$, $g_{2}\coloneqq-A^{\top}x$ are computed. These are the gradients of the utility function of the game. We accelerate the computation by playing the first two actions of the game and then each thread calculates the gradient assuming the first two moves separately and then the gradients are safely reduced. It is worth noting that while there are 81 possible pairs of first moves for the players, only 18 buffers are required to avoid all possible concurrency conflicts. Depending on the machine and the game, calculating the gradients take roughly between 30 and 90 seconds.
@@ A.2 State-of-the-Art Tabular Solvers
    Discounted CFR (DCFR) [Brown and Sandholm, 2019b], in a similar vain as CFR+, discounts the regrets to reduce the effect of previous actions with large (absolute) regret. Practically speaking, at iteration $t$, we rescale the negative regrets by $1-1/(1+t^{\beta})$ and the positive regrets by $1-1/(1+t^{\alpha})$. We use the default value of $\alpha=1.5$ and $\beta=0$.
@@
These are online, first-order optimization algorithms that tabularly refine strategies—represented in sequence-form—by taking steps in the direction of the gradient utility. Table 2 reports the exploitability of equilibrium computed by the methods across the four games and Figure 4 shows the convergence of DCFR and PCFR+. We note that Dark Hex has a winning deterministic strategy for the first moving player. We analyze this strategy in Appendix B.
Table 2: Exploitability after 512 iterations of different algorithms
Game Algorithm Value Exploitability
CFR 0.9999993 0.0000004
CFR+ 0.9999993 0.0000004
DCFR 0.9999993 0.0000004
FP 0.9999997 0.0000002
PCFR 0.9999993 0.0000005
PCFR+ 0.9999993 0.0000005
3x3 Dark Hex (DH3) PDCFR 0.9999993 0.0000005
CFR 0.3841795 0.0295619
CFR+ 0.3844054 0.0011629
DCFR 0.3844321 0.0010063
FP 0.3580483 0.1929647
PCFR 0.3844107 0.0011789
PCFR+ 0.3843936 0.0006123
3x3 Abrupt Dark Hex (ADH3) PDCFR 0.3844073 0.0192829
CFR 0.6664844 0.0107796
CFR+ 0.6666324 0.0016106
DCFR 0.6666511 0.0004070
FP 0.6622969 0.1016475
PCFR 0.6660216 0.0011638
PCFR+ 0.6662418 0.0019337
Phantom Tic-Tac-Toe (PTTT) PDCFR 0.6662910 0.0157007
CFR 0.5501850 0.0162306
CFR+ 0.5501825 0.0004936
DCFR 0.5501728 0.0003577
FP 0.5428041 0.2871734
PCFR 0.5501960 0.0011708
PCFR+ 0.5501886 0.0004846
Abrupt Phantom Tic-Tac-Toe (APTTT) PDCFR 0.5502121 0.0040427
@@ Appendix F Additional results
The evolution of exploitability over 10M training steps for all 4200 hyperparameter tuning runs is shown in Figure 8. Each algorithm-game pair, includes 150 runs – 50 hyperparameter sets with 3 seeds each. The variance across these 3 seeds per set is visualized in Figure 9. Furthermore, we compute the approximate importance of each hyperparameter in the tuning process, which we report in Figure 10. Given the high importance we obtained for the entropy coefficient in the policy-gradient algorithms, in Figure 12 we specifically analyze the average exploitability as a function of the entropy coefficient. We observe that on average we obtain the best exploitability results for entropy coefficients much higher than standard ones detailed in Table 3. Finally, Figure 11 reports the training durations and memory usage of the evaluated algorithms. Note that these metrics depend on hardware and implementation optimizations and are provided for reference only.
Figure 8: Exploitability vs. step for all 4200 runs of the hyperparameter tuning launch, broken down by game and algorithm.
Figure 9: Exploitability for all 4200 runs of the hyperparameter tuning launch, broken down by game and algorithm, then grouped by set of hyperparameters. The boxes-and-whiskers each show the variance over 3 seeds for one set of hyperparameters, and are ordered by decreasing average exploitability.
@@
We show the training exploitability curves for the 1400 runs of the evaluation launch in Figure 13. The variance across seeds for each hyperparameter set can be seen in Figure 9. Additionally, Figure 15 reports the training exploitability curves for the best hyperparameter set of each algorithm-game pair as well as the variance over 10 seeds for each set.
Figure 13: Exploitability vs. step for all 1400 runs of the evaluation launch, broken down by game and algorithm.
Figure 14: Exploitability for all 1400 runs of the evaluation launch, broken down by game and algorithm, then grouped by set of hyperparameters. The boxes-and-whiskers each show the variance over 10 seeds for one set of hyperparameters, and are ordered by decreasing average exploitability.
@@ G.3 R-NaD
    • Description: Schedule for updating the regularization network.
    • Default value: $50{,}000$
    • Notes: This is different from the value of $20{,}000$ in the R-NaD default config but is in the range of our hyperparameter sweep.
@@ G.5 PPO
    • Description: Optimizer learning rate.
    • Default value: $2.5\times 10^{-4}$

### v3

@@ 5 Experiments
While our high-performance exploitability computations open the door to evaluating model-free DRL for IIGs more rigorously, the novelty of these benchmarks necessitates particular care in experimental design and interpretation. Simply showing that tuned PG methods perform “well” on these benchmarks would not suffice as evidence for the policy gradient hypothesis, as there are no numbers for FP-, DO-, or CFR-based DRL algorithms against which to compare. The obvious recourse—which we take—is to report our own results for FP, DO, and CFR-based DRL algorithms. However, since we have hypothesized these algorithms will underperform, our role in tuning them should be carefully scrutinized. To aid readers in this endeavor, we provide detailed documentation of our algorithm implementations, hyperparameter tuning procedure, and hyperparameter tuning results, in addition to final results.
@@ 5.1 Algorithms and Implementation
Figure 2: Exploitability results. For each combination of game and algorithm, the box-and-whisker pair depicts the distribution of final exploitability over the runs from the hyperparameter tuning launch (left) and evaluation launch (right) with square-root y-axis scale. R-NaD, NFSP, ESCHER, and PSRO failed to outperform generic PG methods (MMD, PPO, PPG).
@@ 5.2 Design
For the hyperparameter tuning launch, for each combination of the 5 games and 7 algorithms, we test 50 hyperparameter configurations for 10 million steps over 3 seeds. To create these configurations, we modify the default positive real-valued hyperparameters by independent randomly sampled powers of 2—multiplying for the vast majority of hyperparameters, but exponentiating for some hyperparameters upper bounded by 1. We take the default values for hyperparameters from the implementation source in almost all cases. The most notable exceptions to this are the network architecture and optimizer, for which we impose a 3-layer 512-hidden unit fully connected network and Adam (Kingma and Ba, 2015) on every algorithm, and the entropy coefficient, for which we apply the value from Sokota et al. (2023) to all generic PG methods. We provide a comprehensive list of hyperparameters, default values, and links to the specific lines of code of the external sources from which those default values were taken in Appendix G.
For the evaluation launch, we run the 5 strongest hyperparameter configurations from the first launch (measured by lowest average final exploitability), with 10 fresh seeds, for 10 million steps, for each combination of the 5 games and 7 algorithms. From these evaluation runs, we report both exploitabilities and head-to-head comparisons. For the head-to-head comparisons, we select the policy with the lower median final exploitability over the 10 seeds for each hyperparameter configuration, as well as an approximate Nash equilibrium policy computed using discounted CFR (Brown and Sandholm, 2019b), and compare these selected policies for each pair of algorithms for each game.
@@ 5.3 Results
We summarize two of the main results: exploitability performance and head-to-head performance.
@@
Figure 3 plots head-to-head comparisons for each pair of algorithms. Relative performance patterns are similar to those for exploitability: The generic PG methods matched or defeated other algorithms, while maintaining rough parity among themselves; R-NaD either approached the performance of the generic PG methods or underperformed them, while NFSP and, especially, ESCHER and PSRO were noncompetitive.
@@ 6 Discussion
However, we emphasize that we make no claim—and should not be interpreted as having claimed—that we have established as true the policy gradient hypothesis or any other speculation discussed in this work. A prudent interpretation of our experiments gives consideration to multiple limitations.
First, we only performed experiments on five games, four of which are somewhat homogeneous. We do not believe that the relative performance of the algorithms we evaluated would be universally consistent across adversarial IIGs at large (indeed, it is not even consistent across these five games).
@@ 7 Conclusion
This work hypothesizes and presents evidence that generic deep PG methods are strong approaches for adversarial IIGs. Our results, along with those of Yu et al. (2022), who demonstrated PPO’s effectiveness in cooperative IIGs, and Sokota et al. (2023), who demonstrated MMD’s effectiveness in adversarial IIGs, add to a growing body of evidence for the possibility that a single, simple PG method could serve as a universal algorithm for DRL in games.
@@ Appendix A The exp-a-spiel Package
We release a Python library to efficiently compute exploitability of policies for the five benchmark games introduced in Section 4. The library supports outputting state representations compatible with OpenSpiel [Lanctot et al., 2019], making it straightforward to evaluate the exploitability of policies trained using the latter library.
@@ A.1 Exploitability Computation
Internally, exp-a-spiel evaluates the exploitability of policies by carrying out computations using the so-called sequence-form representation of the five games [Romanovskii, 1962, von Stengel, 1996, Koller et al., 1996]. This representation is roughly 1000 times more compact than the actual game trees, allowing us to compute exact exploitability (and head-to-head values) on commodity hardware. This representation is comprised of two objects: (i) the treeplexes $\cal X,\cal Y$ of the two players [Hoda et al., 2010], and (ii) the sequence-form payoff matrix $A$ of the game. Every policy for player 1 admits an equivalent vector $x\in\cal X$ (resp. $y\in\cal Y$ for player 2), and the expected payoff corresponding to player 1 can be computed in closed form via the bilinear form $x^{\top}Ay$. Due to the size of the games, the sequence-form payoff matrix $A$ is not stored explicitly in memory by exp-a-spiel, but is rather materialized on the fly in a multi-threaded fashion as needed.
@@
    Then, the vectors $g_{1}\coloneqq Ay$, $g_{2}\coloneqq-A^{\top}x$ are computed. These are the gradients of the utility function of the game. We accelerate the computation by playing the first two actions of the game and then each thread calculates the gradient assuming the first two moves separately and then the gradients are safely reduced. It is worth noting that while there are 81 possible pairs of first moves for the players, only 18 buffers are required to avoid all possible concurrency conflicts. Depending on the machine and the game, calculating the gradients take roughly between 30 and 90 seconds. The compute time for calculating best response on various games is presented in Table 2.
@@
Table 2: Runtime for computing best response with exp-a-spiel and OpenSpiel in seconds. ‘OOM’ indicates that the process ran out of memory before timing out. All results in this table were run on an AMD EPYC 7413 24-Core Processor with 256 GB of memory.
Game OpenSpiel exp-a-spiel
Liar’s Dice 1d4f 1 0.03
Liar’s Dice 1d5f 1 0.03
Liar’s Dice 1d6f 5 0.04
Liar’s Dice 2d3f 10 0.04
Liar’s Dice 2d4f 527 0.45
Liar’s Dice 2d5f OOM 11
Classical PTTT OOM 27
Abrupt PTTT OOM 27
Classical DH3 OOM 47
Abrupt DH3 OOM 47
@@ A.2 State-of-the-Art Tabular Solvers
    Discounted CFR (DCFR) [Brown and Sandholm, 2019a], in a similar vain as CFR+, discounts the regrets to reduce the effect of previous actions with large (absolute) regret. Practically speaking, at iteration $t$, we rescale the negative regrets by $1-1/(1+t^{\beta})$ and the positive regrets by $1-1/(1+t^{\alpha})$. We use the default value of $\alpha=1.5$ and $\beta=0$.
@@
These are online, first-order optimization algorithms that tabularly refine strategies—represented in sequence-form—by taking steps in the direction of the gradient utility. Table 3 reports the exploitability of equilibrium computed by the methods across the 3x3 games and Figure 4 shows the convergence of DCFR and PCFR+. We note that Dark Hex has a winning deterministic strategy for the first moving player. We analyze this strategy in Appendix B.
Table 3: Exploitability after 512 iterations of different algorithms on 3x3 board games
[…]
@@ D.2 MMD
Table 4: Related work and their use of exact or approximate exploitability when evaluating their method.
Paper Exact Exploitability Approx. Exploitability or Head-to-Head
NFSP [Heinrich and Silver, 2016] Leduc
PSRO [Lanctot et al., 2017] Leduc
ESCHER [McAleer et al., 2023] 2x2 Battle Ship!, Leduc Dark Chess, DH5, PTTT
RNaD [Perolat et al., 2022] Stratego
MMD [Sokota et al., 2023] LD1D4F, Leduc, ADH2 ADH3, PTTT
Qin et al. [2019] Kuhn Poker
AdaptFSP [Goldstein and Brown, 2023] Leduc
Hu et al. [2023] Leduc
NSFP-PLT [Li et al., 2023a] 6-, 3-, 2-player poker, Leduc
Pipeline PSRO [McAleer et al., 2020] Leduc Barrage
XDO [McAleer et al., 2021] Leduc
Smith et al. [2021] Leduc
Anytime PSRO[McAleer et al., 2022b] Leduc
Online DO [Dinh et al., 2022]
Zhou et al. [2022] Leduc
PDO [Yao et al., 2023] Leduc
SP-PSRO [McAleer et al., 2022a] LD1D6F, 2x2 Battle Ship!, Leduc
Self-Adaptive PSRO [Li et al., 2024b] LD1D6F, Leduc
Fusion PSRO [Lian, 2024] LD1D6F, Leduc
Dream [Steinberger et al., 2020] Leduc Flop Hold’em
ARMAC [Gruslys et al., 2020] LD1D6F, Leduc Head’s up
FoReL [Perolat et al., 2021] LD1D6F, Leduc
Meng et al. [2023] Leduc PTTT, DH5, DH4
Liu et al. [2024] Leduc
Table 5: Size of games for which exact exploitability is reported in prior work.
Game Number of Infosets Number of States (including terminal)
ADH2 94 471
Leduc 936 9457
LD1D4F 1024 $81$
2x2 Battle Ship! 3286 10069
LD1D6F 24576 294883
@@ Appendix F Additional results
The evolution of exploitability over 10M training steps for all 5250 hyperparameter tuning runs is shown in Figure 8. Each algorithm-game pair, includes 150 runs—50 hyperparameter sets with 3 seeds each. The variance across these 3 seeds per set is visualized in Figure 9. Furthermore, we compute the approximate importance of each hyperparameter in the tuning process, which we report in Figure 10. Given the high importance we obtained for the entropy coefficient in the policy-gradient algorithms, in Figure 12 we specifically analyze the average exploitability as a function of the entropy coefficient. We observe that on average we obtain the best exploitability results for entropy coefficients much higher than standard ones detailed in Table 6. We also performed some additional experiments on a wider hyperparameter sampling range for the number_training_episodes hyperparameter in PSRO, shown in Table 7. Figure 11 reports the training durations and memory usage of the evaluated algorithms. Note that these metrics depend on hardware and implementation optimizations and are provided for reference only.
Table 7: This table shows the performance of PSRO on our benchmark games using a wider sampling range for the number_training_episodes hyperparameter that determines for how long each oracle trains. Exploitability is presented as percentiles across $50$ different hyperparameter samples. Results for the original sampling range presented in the main results of the paper ($[125,8000]$) are shown for comparison
Game Sampling Range expl. @ 10% expl. @ 25% expl. @ 50% expl. @ 75%
ADH3 $[2500,160000]$ 0.929014 1 1 1
ADH3 $[125,8000]$ 0.750014 0.894206 1 1
APTTT $[2500,160000]$ 0.934627 1 1 1
APTTT $[125,8000]$ 0.795704 0.893034 0.988603 1
DH3 $[2500,160000]$ 0.969553 1 1 1
DH3 $[125,8000]$ 0.72258 0.862278 1 1
PTTT $[2500,160000]$ 0.88345 1 1 1
PTTT $[125,8000]$ 0.786203 0.886317 0.99254 1
Figure 8: Exploitability vs. step for all 5250 runs of the hyperparameter tuning launch, broken down by game and algorithm.
Figure 9: Exploitability for all 5250 runs of the hyperparameter tuning launch, broken down by game and algorithm, then grouped by set of hyperparameters. The boxes-and-whiskers each show the variance over 3 seeds for one set of hyperparameters, and are ordered by decreasing average exploitability.
@@
We show the training exploitability curves for the 1750 runs of the evaluation launch in Figure 13. The variance across seeds for each hyperparameter set can be seen in Figure 9. Additionally, Figure 15 reports the training exploitability curves for the best hyperparameter set of each algorithm-game pair as well as the variance over 10 seeds for each set.
Figure 13: Exploitability vs. step for all 1750 runs of the evaluation launch, broken down by game and algorithm.
Figure 14: Exploitability for all 1750 runs of the evaluation launch, broken down by game and algorithm, then grouped by set of hyperparameters. The boxes-and-whiskers each show the variance over 10 seeds for one set of hyperparameters, and are ordered by decreasing average exploitability.
@@ G.3 R-NaD
    • Description: Number of samples in each training batch.
    • Default value: $256$
@@
    • Description: Coefficient for V-trace importance weights [Espeholt et al., 2018].
    • Default value: $1.0$